1. integrate |1-x^2|dx from limit -2 to 2.
2. Integrate |2x+3|dx from limit -2 to 2
1
Expert's answer
2020-05-15T17:26:57-0400

The problem (1):-


Since "1-x^{2}>0" when "-1<x<1" and then "\left|1-x^{2}\right|=1-x^{2}"

"\begin{array}{l}
\text { But } 1-x^{2}<0 \quad \text { when } \quad 1<x \text { or } x<-1, \text { thus }

 \\
\left|1-x^{2}\right|=x^{2}-1
\end{array}"

Thus

"\begin{array}{l}
\int_{-2}^{2}\left|1-x^{2}\right| d x \\
=\int_{-2}^{-1}\left|1-x^{2}\right| d x+\int_{-1}^{1}\left|1-x^{2}\right| d x+\int_{1}^{2}\left|1-x^{2}\right| d x \\
=\int_{-2}^{-1}\left(x^{2}-1\right) d x \\
+\int_{-1}^{1}\left(1-x^{2}\right) d x+\int_{1}^{2}\left(x^{2}-1\right) d x \\
=\left(\frac{x^{3}}{3}-x\right)_{-2}^{-1}+\left(x-\frac{x^{3}}{3}\right)_{-1}^{1}+\left(\frac{x^{3}}{3}-x\right)_{1}^{2} \\
=\left(\frac{-1}{3}+1\right)-\left(\frac{-8}{3}+2\right)+\left(\frac{-1}{3}+1\right) \\
=4
\end{array}"


The problem (2):-

Since "2 x+3>0" when "x>\frac{-3}{2}" , then "|2 x+3|=2 x+3"

But "2x+3<0" when "x<\frac{-3}{2}" , thus "|2 x+3|=-2 x-3"

Therefore we can find the value of the following integral

"\begin{array}{l}
\int_{-2}^{2}|2 x+3| d x=\int_{-2}^{\frac{-3}{2}}|2 x+3| d x
+\int_{-3}^{2}|2 x+3| d x \\
=\int_{-2}^{\frac{-3}{2}}(-2 x-3) d x+\int_{\frac{-3}{2}}^{2}(2 x+3) d x \\
=\left(-x^{2}-3 x\right)_{-2}^{\frac{-3}{2}}+\left(x^{2}+3 x\right)_{\frac{-3}{2}}^{2} \\

=\left(-\frac{9}{4}+\frac{9}{2}\right)-(-4+6)+(4+6)-\left(\frac{9}{4}-\frac{9}{2}\right) \\
=\frac{25}{2}
\end{array}"